OpenStudy (anonymous):

PLEASE HELP ASAP!! what is the equation of a circle if, Center (3, -2), radius = 5

OpenStudy (anonymous):

(x-h)^2+(y-k)^=r^2

OpenStudy (anonymous):

this is for a circle with c(h,k)

OpenStudy (campbell_st):

the general form of a circle is \[(x -h)^2 + (y - k)^2 = r^2\] with centre at (h, k) and radius r in your question you have centre (3, -2) and r = 5 just substitute

OpenStudy (anonymous):

now you can substitue

OpenStudy (anonymous):

thank you guys!!!!